the solution set of the equation x^2-4x=0 is?
Answer by tutorcecilia(2152)

x^2-4x=0 [factor whereever possible]
x(x-4)=0 [set each factor equal to zero and solve for x]
x=0
and
x-4=0
x=4
.
check by plugging each factor back into the orginal equation and solve.
x^2-4x=0
Let x=0
0^2=4(0)=0
0=0 [checks out]
and
x^2-4x=0
(4)^2-4(4)x=0
16-16=0
0=0 [also checks out]